OpenStudy (anonymous):

You are finding the product of two square roots. Each radicand is the same variable raised to an odd power. Explain why you would want to multiply the radicands together first and then simplify the product.

OpenStudy (anonymous):

(x)^3 * (x)^5 ....... well why do you think so ;)

OpenStudy (anonymous):

because if you add two odd numbers, you will get an even number, and taking the square root will be the same as dividing the exponent by 2

OpenStudy (anonymous):

\[\sqrt{x^5}\sqrt{x^7}=\sqrt{x^{12}}=x^6\]for example